OpenStudy (leann01xxx):

Select a counter-example that makes the conclusion false. 7 - 3 = 4, 8 - 5 = 3, 9 - 8 = 1 Conclusion: the difference of two positive numbers is always positive. Counterexample:

OpenStudy (leann01xxx):

I have to write out how to explain how to get the answer and i dont know how to do either.. :(

rishavraj (rishavraj):

7 - 9 = -2

OpenStudy (leann01xxx):

I have to explain how to get the answer

rishavraj (rishavraj):

hmmm wht answer ??

OpenStudy (leann01xxx):

I dont know or an example or something

rishavraj (rishavraj):

like see in the example 8 - 1 = 7 first number is larger .,.so the answer is positive but if the second number was larger then the answer would be negative

OpenStudy (leann01xxx):

Okay.. so, now what?

rishavraj (rishavraj):

see if the second number is greater like 2 - 30 the easiest thing to do is to reverse the numbers and then make the answer negative after you've subtracted them. So, 30 - 2 = 28, since 28 is just two less than 30. Now, make your answer negative since you determined at the beginning that it would be negative because the second number is larger than the first. Therefore, 2 - 30 = -28.

OpenStudy (leann01xxx):

Im sorry i just dont know how to do this

OpenStudy (leann01xxx):

So is that what i would write down then?

OpenStudy (leann01xxx):

@rishavraj

rishavraj (rishavraj):

@LeAnn01Xxx I think so ...... :/

OpenStudy (leann01xxx):

Alright. I'll just put it and if its wrong oh well then, thanks tho.

rishavraj (rishavraj):

yeah just go with it...... and mention examples too

OpenStudy (leann01xxx):

Okay

rishavraj (rishavraj):

and u welcome :)) and I really hope its correct :P